OpenStudy (anonymous):

A marble is drawn from a bag containing 3 red marbles and 2 blue marbles and a number cube with faces numbered 1 through 6 is tossed. What is the probability that the marble drawn is blue and the number cube shows a number less than 3? A marble is drawn from a bag containing 3 red marbles and 2 blue marbles and a number cube with faces numbered 1 through 6 is tossed. What is the probability that the marble drawn is blue A.11/15 B.1/4 C.1/6 D.2/15

OpenStudy (anonymous):

The probability that he will draw a blue marble \[ \frac{ \text{ blue marbles }}{\text{ total amount of marbles }}=\frac{2}{5}\]

OpenStudy (anonymous):

The probability that the die will land on a number less than 3 namely 1 or 2 is \[ \frac{2}{6}\]

OpenStudy (anonymous):

2/5 isnt in here though

OpenStudy (anonymous):

So the probability that both will occur is (2/5)*(2/6)

OpenStudy (anonymous):

its 1/6

OpenStudy (anonymous):

Ummm use you calculator if you cant multiply fractions -.-

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

\[\frac{2}{5}*\frac{2}{6}=\frac{4}{30}=\frac{2}{15}\]

OpenStudy (anonymous):

Anyways Im out